# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ 4.2. Explicit versions of theorems of Győry (1973) and Evertse and Győry (1991a)
5. Consequences of Theorem 4.2 in algebraic number theory, and in particular for monogenicity of number fields and their orders
@@
Theorem 4.2 and its variants in Győry (1973, 74, 76, 78a,b) led to breakthroughs in the effective theory of number fields. They furnished general effective finiteness results for integral elements of given non-zero discriminant resp. given index in any number field $K$ and its orders. In particular, in Győry (1976), Theorem 4.2 provided the first general algorithm for deciding the monogenicity of $K$ and, more generally, of any order of $K$, and for determining all power integral bases in $K$ and in its orders. For later generalizations, applications and comprehensive treatment of this extensive area, we refer to Győry (1980a, 1980b, 2000), Evertse and Győry (2017), the references given there, and to Section 8 of the present paper.
For convenience, we present the general effective finiteness theorems mentioned above in their simplest form. Further, we briefly sketch how to deduce the consequences directly from Theorem 4.2 via its Corollary 5.1.
@@ 5.1. Preliminaries
Throughout this section, $K$ will denote a number field of degree $n\geq 2$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$. For a primitive integral element $\alpha$ of $K$ we denote by $f_{\alpha}(X)$ the minimal (monic) polynomial of $\alpha$ in $\mathbb{Z}[X]$. Thus, $f_{\alpha}(X)=\prod_{i=1}^{n}(X-\alpha^{(i)})$, where $n$ is the degree of $\alpha$ and $\alpha^{(1)}=\alpha,\alpha^{(2)},\ldots,\alpha^{(n)}$ denote the conjugates of $\alpha$. Then the discriminant of $\alpha$ relative to the extension $K/\mathbb{Q}$ is defined by
(5.1) $D_{K/\mathbb{Q}}(\alpha):=D(f_{\alpha})=\prod_{1\leq i<j\leq n}(\alpha^{(i)}-\alpha^{(j)})^{2}.$
Let now $\mathcal{O}$ be an order of $K$ (i.e., a subring of $K$ that as a $\mathbb{Z}$-module is free of rank $n=[K:\mathbb{Q}]$), and $D_{\mathcal{O}}$ its discriminant. Then $\mathcal{O}$ is a subring of $\mathcal{O}_{K}$. For $\alpha\in\mathcal{O}_{K}$ resp. $\alpha\in\mathcal{O}$, we denote by $I(\alpha):=[\mathcal{O}_{K}:\mathbb{Z}[\alpha]]$ resp. by $I_{\mathcal{O}}(\alpha):=[\mathcal{O}:\mathbb{Z}[\alpha]]$ the index of $\alpha$ in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$. Then, as is known,
(5.2a) $D_{K/\mathbb{Q}}(\alpha)=I^{2}(\alpha)D_{K}\text{ if }\alpha\in\mathcal{O}_{K},$
and
(5.2b) $D_{K/\mathbb{Q}}(\alpha)=I^{2}_{\mathcal{O}}(\alpha)D_{\mathcal{O}}\text{ if }\alpha\in\mathcal{O},$
where $D_{\mathcal{O}}$ is the discriminant of $\mathcal{O}$, i.e., of any $\mathbb{Z}$-module basis of $\mathcal{O}$.
@@
Clearly, $\mathbb{Z}$-equivalence elements in $\mathcal{O}_{K}$ resp. $\mathcal{O}$ have the same discriminant and hence the same index in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$.
@@
Further, $K$ resp. $\mathcal{O}$ is called $k\,(\geq 1)$ times monogenic if $\mathcal{O}_{K}$ resp. $\mathcal{O}$ equals $\mathbb{Z}[\alpha_{1}]=\cdots=\mathbb{Z}[\alpha_{k}]$ for some pairwise $\mathbb{Z}$-inequivalent $\alpha_{1},\ldots,\alpha_{k}$ in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$. In case that in the above definition $k$ is maximal, it is called the multiplicity of the monogenicity of $K$, resp. $\mathcal{O}$.
5.2. Most important consequences of Theorem 4.2 in number fields
@@
By the height $H(\alpha)$ of an algebraic integer $\alpha$ we mean the height $H(f_{\alpha})$. Further, the discriminant $D(\alpha)$ of $\alpha$ is defined as $D(\alpha):=D(f_{\alpha})$.
Corollary 5.1 (of Theorem 4.2).
@@
  2. (ii)
    there is an algebraic integer $\beta$, $\mathbb{Z}$-equivalent to $\alpha$, such that
@@
This implies that there are only finitely many $\mathbb{Z}$-equivalence classes of algebraic integers with a given non-zero discriminant, and that a full set of representatives of these classes can be effectively determined.
Corollary 5.1 was proved in Győry (1973) as ‘Corollaire 3’ of the ‘Théorème’, the main result of the paper. This finiteness result in an ineffective form follows also from the work of Birch and Merriman (1972), which was independent of Győry (1973). Corollary 5.1 confirmed in full generality and in effective form a conjecture of Nagell (1967). Previously, the cubic case was settled independently by Delone (1930) and Nagell (1930), and the quartic case by Nagell (1967).
Finally, we note that Corollary 5.1 easily follows from Theorem 4.2. Indeed, if $\alpha$ is an algebraic integer with the properties specified in Corollary 5.1, then by (5.1), $D(f_{\alpha})=D$ and $\deg f_{\alpha}=n$. Further, by Theorem 4.2 $f_{\alpha}$ is $\mathbb{Z}$-equivalent to some monic $g\in\mathbb{Z}[X]$ with degree $n$ and discriminant $D$ such that $n\leq c_{1}(|D|)$ and $H(g)\leq c_{2}(n,|D|)$, where $c_{1},c_{2}$ denote the effectively computable numbers occurring in Theorem 4.2. But then $\alpha$ is $\mathbb{Z}$-equivalent to a zero of $g$, say $\beta$, whence $\deg\beta\leq c_{1}(|D|)$ and $H(\beta)\leq c_{2}(n,|D|)$ follow.∎
We note that Corollary 5.1 above and also Corollaire 1 of Győry (1973) imply in an effective way that there are only finitely many algebraic units in $\overline{\mathbb{Q}}$ with given discriminant. This gave the effective solution to Problem 19 in the book Narkiewicz (1974).
Consider again the above number field $K$ of degree $n$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$, and let $\mathcal{O}$ be an order of $K$ with discriminant $D_{\mathcal{O}}$. Restricting ourselves in Corollary 5.1 to the integers in $\mathcal{O}$, we get the following.
@@
Let $\mathcal{O}$ be an order of $K$, and let $\alpha\in\mathcal{O}$ be a primitive element of $K$ with index $I_{\mathcal{O}}$ in $\mathcal{O}$. Then $\alpha$ is $\mathbb{Z}$-equivalent to some $\beta$ in $\mathcal{O}$ such that
$H(\beta)\leq c_{2}(n,I^{2}_{\mathcal{O}}\cdot|D_{\mathcal{O}}|).$
If in particular $\mathcal{O}=\mathcal{O}_{K}$ and $\alpha\in\mathcal{O}_{K}$ with index $I$ in $\mathcal{O}_{K}$, then the same holds with $I_{\mathcal{O}},\,D_{\mathcal{O}}$ replaced by $I$ and $D_{K}$ respectively, where $c_{2}$ denotes the effectively computable number occurring in Theorem 4.2.
This implies that there are only finitely many $\mathbb{Z}$-equivalence classes of $\alpha$ in $\mathcal{O}$ resp. in $\mathcal{O}_{K}$ with given index $I_{\mathcal{O}}$ in $\mathcal{O}$ resp. $I$ in $\mathcal{O}_{K}$, and a full set of representatives of these classes can be effectively determined.
Corollary 5.2 follows immediately from Corollary 5.1, using (5.2b) resp. (5.2a). It was proved in Győry (1976) in an effective and quantitative form.
Corollary 5.3 below is the most influential consequence of Theorem 4.2. It gave the first general algorithm for deciding the monogenicity, and multiplicity of monogenicity of $K$ and its orders $\mathcal{O}$, and for determining all power integral bases in $K$ and its orders.
@@
If $\alpha$ is an element of $\mathcal{O}$ resp. of $\mathcal{O}_{K}$ such that $\mathcal{O}=\mathbb{Z}[\alpha]$ resp. $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ then $\alpha$ is $\mathbb{Z}$-equivalent to some $\beta$ in $\mathcal{O}$ resp. in $\mathcal{O}_{K}$ such that its height $H(\beta)$ is at most $c_{2}(n,|D_{\mathcal{O}}|)$ resp. $c_{2}(n,|D_{K}|)$, where $c_{2}$ denotes the effectively computable expression in Theorem 4.2 with $|D|$ replaced by $|D_{\mathcal{O}}|$ resp. by $|D_{K}|$.
It follows from Corollary 5.3 that there are only finitely many $\mathbb{Z}$-equivalence classes of $\alpha$ in $\mathcal{O}$ resp. in $\mathcal{O}_{K}$ such that $\mathcal{O}=\mathbb{Z}[\alpha]$ resp. $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$, and a full set of representatives of these classes can be effectively determined.
Corollary 5.3 is an immediate consequence of Corollary 5.2, choosing $I_{\mathcal{O}}=1$ resp. $I=1$. It was established in Győry (1976) with explicit expressions for $c_{1}$ and $c_{2}$.
@@
With the above formulation of Corollaries 5.1 to 5.3 it was easier to point out that these corollaries are indeed consequences of Theorem 4.2. Further, we note that explicit versions of the corollaries can be easily derived from the explicit variant Theorem 4.2* of Theorem 4.2. Finally, the corollaries can be deduced with better bounds from less general versions of Theorem 4.2, where the polynomials $f$ involved are irreducible; for such versions we refer to Győry (1976, 1998, 2000), Evertse and Győry (2017) and in fact Corollary 5.1 above.
5.3. Reformulation of Corollaries 5.1 (ii), 5.2 and 5.3 in terms of polynomial Diophantine equations over $\mathbb{Z}$
@@
Let $K$ be an algebraic number field of degree $n\geq 2$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$. Let $\{1,\omega_{2},\ldots,\omega_{n}\}$ be an integral basis of $K$. For $\alpha\in\mathcal{O}_{K}$ with
@@
can be regarded as a decomposable form of degree $n(n-1)$ in $x_{2},\ldots,x_{n}$ with coefficients in $\mathbb{Z}$, i.e., it is a product of $n(n-1)$ linear forms in $x_{1},\ldots,x_{n}$ with algebraic coefficients. The form $D(x_{2}\omega_{2}+\cdots+x_{n}\omega_{n})$, which was introduced by Kronecker (1882), is called discriminant form, while, for $D\neq 0$, the equation
@@
Clearly, Corollary 5.1, (ii) is equivalent to the following
@@
to the integral basis $\{1,\omega_{2},\ldots,\omega_{n}\}$ of $K$ there corresponds a decomposable form $I(X_{2},\ldots,X_{n})$ of degree $n(n-1)/2$ in $n-1$ variables with coefficients in $\mathbb{Z}$ such that for $\alpha\in\mathcal{O}_{K}$
@@
We note that the equations (5.4) and (5.6) are related by (5.2a).
In view of (5.5), the finiteness assertion of Corollary 5.2 for $\mathcal{O}_{K}$ is equivalent to the following.
@@
In particular, for $I=1$, we get the following equivalent formulation of Corollary 5.3 for $\mathcal{O}_{K}$.

### v2

@@ 5. Consequences in algebraic number theory, in particular for monogenicity and rational monogenicity
We give some consequences of Theorems 4.1, 4.2 and 4.4 in algebraic number theory. Of particular interest are applications to monogenicity of number fields and (rational) monogenicity of orders.
Theorem 4.1 due to Birch and Merriman from 1972 has an important ineffective finiteness consequence for algebraic integers of given discriminant; see Theorem 5.1 below.
An effective version of Theorem 5.1 was obtained independently in Győry (1973), as a consequence of his effective Theorem 4.2 presented above; see Theorem 5.2 below.
@@
Theorem 5.2 as well as its various effective consequences, applications, quantitative variants and generalizations in Győry (1973, 1974, 1976, 1978a,b, 1980a,b, 1981) led to breakthroughs in the effective theory of number fields. These furnished among others general effective finiteness results for integral elements of given discriminant resp. of given index in number fields and, more generally, in their orders; see Corollaries 5.3 and 5.4. In particular, as an immediate consequence of his Theorem 5.2, Győry provided the first general effective algorithm for deciding the monogenicity and for determining, at least in principle, all power integral bases in number fields and in their orders; see Theorems 5.5 and 5.6 below.
As a consequence of Theorem 4.4 we present from Evertse and Győry (1991a) a general effective finiteness theorem on algebraic numbers of given discriminant; see Theorem 5.10. Finally, we introduce rationally monogenic orders, which are generalizations of monogenic orders, and give an algorithm to determine in principle whether a given order is rationally monogenic, see Theorems 5.14 and 5.15 below.
For convenience, we formulate the above-mentioned effective finiteness results in their simplest form. For generalizations, further applications and comprehensive treatment of this extensive area, we refer to Győry (1983, 1984, 1998, 2000, 2006), Evertse and Győry (1991a, 2017, 2022), BEGyRS (2023), the references given there, and to Sections 6 to 9 of the present paper.
@@ 5.1. Preliminaries
Throughout this section, $K$ will denote a number field of degree $n\geq 2$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$. Recall that $K$ has precisely $n$ distinct embeddings in its normal closure over $\mathbb{Q}$, which we denote by $x\mapsto x^{(i)}$ ($i=1,\ldots,n$). Here $x^{(1)}=x$.
Let $\mathcal{M}$ be a free $\mathbb{Z}$-module in $K$ of rank $n$, and pick a $\mathbb{Z}$-basis $\{\omega_{1},\ldots,\omega_{n}\}$ of $\mathcal{M}$. Then the discriminant of $\mathcal{M}$ is defined by
$D(\mathcal{M}):=\Big(\det\big(\omega_{i}^{(j)}\big)_{i,j=1}^{n}\Big)^{2}.$
This is a rational number, and it does not depend on the choice of the basis.
Given two free $\mathbb{Z}$-modules $\mathcal{M}_{1},\mathcal{M}_{2}$ in $K$ of rank $n$ with $\mathcal{M}_{1}\supseteq\mathcal{M}_{2}$, denote by $\mathcal{M}_{1}:\mathcal{M}_{2}]$ the index of $\mathcal{M}_{2}$ in $\mathcal{M}_{1}$, i.e., the cardinality of $\mathcal{M}_{1}/\mathcal{M}_{2}$. Then
(5.1) $D(\mathcal{M}_{2})=[\mathcal{M}_{1}:\mathcal{M}_{2}]^{2}D(\mathcal{M}_{1}).$
We recall that an order of $K$ is a subring of $K$ which as a $\mathbb{Z}$-module is free of rank $n$. In particular, $\mathcal{O}_{K}$ is an order of $K$, and each other order of $K$ is a subring of $\mathcal{O}_{K}$. We have $D_{K}=D(\mathcal{O}_{K})$. Equation (5.1) implies that if $\mathcal{M}$ is a submodule of $\mathcal{O}_{K}$ of rank $n$, then $D(\mathcal{M})\in\mathbb{Z}$.
Let $\alpha$ be a non-zero algebraic integer. Then we denote by $f_{\alpha}(X)$ the minimal (monic) polynomial of $\alpha$ in $\mathbb{Z}[X]$. Thus, $f_{\alpha}(X)=\prod_{i=1}^{n}(X-\alpha^{(i)})$, where $\alpha^{(1)}=\alpha,\ldots,\alpha^{(n)}$ are the distinct conjugates of $\alpha$ in the splitting field of $f_{\alpha}$. We now define the discriminant of $\alpha$ by
(5.2) $D(\alpha):=D(f_{\alpha})=\prod_{1\leq i<j\leq n}(\alpha^{(i)}-\alpha^{(j)})^{2}.$
The ring $\mathbb{Z}[\alpha]$ is clearly an order of $\mathbb{Q}(\alpha)$, with $\mathbb{Z}$-module basis $\{1,\alpha,\ldots,\alpha^{n-1}\}$, so
(5.3) $D(\alpha)=D(\mathbb{Z}[\alpha]).$
Let now $\mathcal{O}$ be an order of $K$, and $D(\mathcal{O})$ its discriminant. Then $\mathcal{O}$ is a subring of $\mathcal{O}_{K}$. For a primitive element $\alpha$ of $K$ with $\alpha\in\mathcal{O}_{K}$ resp. $\alpha\in\mathcal{O}$, we define
(5.4) $I(\alpha):=[\mathcal{O}_{K}:\mathbb{Z}[\alpha]],\ \ I_{\mathcal{O}}(\alpha):=[\mathcal{O}:\mathbb{Z}[\alpha]]$
to be the index of $\alpha$ in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$. Then, by (5.1), (5.3),
(5.5) $D(\alpha)=I(\alpha)^{2}D_{K}\text{ for }\alpha\in\mathcal{O}_{K},\ \ D(\alpha)=I_{\mathcal{O}}(\alpha)^{2}D(\mathcal{O})\text{ for }\alpha\in\mathcal{O}.$
@@
Clearly, $\mathbb{Z}$-equivalent elements in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$ have the same discriminant and hence the same index in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$.
@@
Further, $K$ resp. $\mathcal{O}$ is called $k\,(\geq 1)$ times monogenic if $\mathcal{O}_{K}$ resp. $\mathcal{O}$ equals $\mathbb{Z}[\alpha_{1}]=\cdots=\mathbb{Z}[\alpha_{k}]$ for some pairwise $\mathbb{Z}$-inequivalent $\alpha_{1},\ldots,\alpha_{k}$ in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$. In case that in the above definition $k$ is maximal, it is called the multiplicity of the monogenicity of $K$, resp. of $\mathcal{O}$.
5.2. Consequences of Theorems 4.1 and 4.2 for algebraic integers of given discriminant
@@
From their Theorem 4.1, Birch and Merriman in 1972 deduced the following ineffective finiteness theorem.
Theorem 5.1 (Birch and Merriman, 1972).
Up to $\mathbb{Z}$-equivalence, there are only finitely many algebraic integers with given non-zero discriminant.
Independently, as a consequence of his Theorem 4.2, Győry (1973) proved the following effective version of Theorem 5.1.
By the height $H(\alpha)$ of an algebraic integer $\alpha$ we mean the height $H(f_{\alpha})$.
Theorem 5.2 (Győry, 1973).
@@
  • (ii)
    There is an algebraic integer $\beta$, $\mathbb{Z}$-equivalent to $\alpha$ such that
@@
This theorem was stated and proved in Győry (1973) as ’Corollaire 3’ of the ’Théorème’, cf. Theorem 4.2 above.
As was mentioned in Section 2, the cubic case was settled independently by Delone (1930) and Nagell (1930), and the quartic case by Nagell (1967) in an ineffective way.
Theorems 5.1 resp. 5.2 confirmed in full generality, and in fact Theorem 5.2 in an effective form, a conjecture of Nagell (1967). Further both Theorem 5.1 and Theorem 5.2 imply, Theorem 5.2 in an effective form, that there are only finitely many algebraic units in $\overline{\mathbb{Q}}$ of given discriminant. This gave the effective solution to Problem 19 in the book Narkiewicz (1974).
Finally, we note that Theorem 5.2 easily follows from Theorem 4.2. Indeed, if $\alpha$ is an algebraic integer with the properties specified in Theorem 5.2, then by (5.2), $D(f_{\alpha})=D$ and $\deg f_{\alpha}=n$. Further, by Theorem 4.2 $f_{\alpha}$ is $\mathbb{Z}$-equivalent to some monic $g\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D$ such that $n\leq c_{1}(|D|)$ and $H(g)\leq c_{2}(n,|D|)$, where $c_{1},c_{2}$ denote the effectively computable numbers occurring in Theorem 4.2. But then $\alpha$ is $\mathbb{Z}$-equivalent to a zero of $g$, say $\beta$, whence $\deg\beta\leq c_{1}(|D|)$ and $H(\beta)\leq c_{2}(n,|D|)$ follow.∎
The first explicit version of (ii) in Theorem 5.2 was established by Győry (1974) by means of Baker’s method. For $c_{1}(n)$ one can take $2\log|D|/\log 3$. For $c_{2}(n,|D|)$ we can obtain an explicit bound, using Theorem 4.2* instead of Theorem 4.2. An even better explicit estimate can be obtained in (ii), observing that in fact we apply Theorem 4.2 (or its explicit version Theorem 4.2*) only to irreducible polynomials $f_{\alpha}$. The best known bound in (ii) comes from Theorem 6.4.1 of Evertse and Győry (2017).
5.3. Consequences for monogenic number fields and orders
a
Let again $K$ be a number field of degree $n\geq 2$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$.
The following effective corollaries are immediate consequences of Theorem 5.2 (i.e. the ‘Corollaire 3’) of Győry (1973). Although this was not mentioned by Birch and Merriman in their 1972 paper, it should be remarked that from their Theorem 5.1 one can also deduce in ineffective form the finiteness consequences of the results below.
@@ 5.3. Consequences for monogenic number fields and orders
Let $\mathcal{O}$ be an order of $K$ and $D$ a non-zero integer. Every $\alpha$ in $\mathcal{O}$ of discriminant $D_{K/\mathbb{Q}}(\alpha)=D$ is $\mathbb{Z}$-equivalent to some $\beta\in\mathcal{O}$ such that
$H(\beta)\leq c_{2}(n,|D|),$
where $c_{2}=c_{2}(n,|D|)$ denotes the same effectively computable positive number as in Theorem 5.2.
This is a special case of Theorem 5.2, restricted to the elements of $\mathcal{O}$. It follows from Corollary 5.3 that up to $\mathbb{Z}$-equivalence, there are only finitely many elements of $\mathcal{O}$ of given non-zero discriminant, and all of them can be, at least in principle, effectively determined.
As was mentioned above, the first quantitative versions of Theorem 5.2 and Corollary 5.3 were established in Győry (1974).
@@
Let $\mathcal{O}$ be an order in $K$ of discriminant $D(\mathcal{O})$, and $I_{\mathcal{O}}$ a positive integer. Every $\alpha$ in $\mathcal{O}$ with index $I_{\mathcal{O}}(\alpha)=I_{\mathcal{O}}$ is $\mathbb{Z}$-equivalent to some $\beta\in\mathcal{O}$ such that
$H(\beta)\leq c_{2}(n,I^{2}_{\mathcal{O}}\cdot|D(\mathcal{O})|),$
where $c_{2}$ denotes the same effectively computable positive number as in Theorem 5.2 with $|D|$ replaced by $I^{2}_{\mathcal{O}}\cdot|D(\mathcal{O})|$.
This follows immediately from Corollary 5.3 and the second identity in (5.5). Corollary 5.4 implies that up to $\mathbb{Z}$-equivalence there are only finitely many elements in $\mathcal{O}$ with given index and all of them can be, at least in principle, effectively determined.
The next Theorem 5.5 and its more general version Theorem 5.6 are the most influential consequences of Theorem 5.2. They provided the first general effective algorithm for deciding the monogenicity, the multiplicity of monogenicity, and for determining, at least in principle, all power integral bases in $K$ and in its orders.
Of particular importance are the cases when in Corollaries 5.3, and 5.4 $\mathcal{O}$ is just $\mathcal{O}_{K}$, the ring of integers of $K$. Then Corollary 5.4 implies
Theorem 5.5 (Győry, 1976).
Every $\alpha\in\mathcal{O}_{K}$ with $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ is $\mathbb{Z}$-equivalent to some $\beta\in\mathcal{O}_{K}$ such that
$H(\beta)\leq c_{2}(n,|D_{K}|),$
where $c_{2}$ denotes the same effectively computable positive number as in Corollary 5.4 with $I_{\mathcal{O}}=1$, $D(\mathcal{O})=D_{K}$. Consequently, there are only finitely many $\mathbb{Z}$-equivalence classes of $\alpha$ in $\mathcal{O}_{K}$ such that $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$, and a full set of representatives of these classes can be, at least in principle, effectively found.
More generally, Corollary 5.4 immediately gives the following:
Theorem 5.6 (Győry, 1976).
Let $\mathcal{O}$ be an order of $K$ of discriminant $D(\mathcal{O})$. Every $\alpha\in\mathcal{O}$ with $\mathcal{O}=\mathbb{Z}[\alpha]$ is $\mathbb{Z}$-equivalent to some $\beta\in\mathcal{O}$ such that
$H(\beta)\leq c_{2}(n,|D(\mathcal{O})|),$
where $c_{2}$ denotes the same effectively computable positive number as in Corollary 5.4 with $I_{\mathcal{O}}=1$.
The first explicit, quantitative versions of Corollary 5.4 and Theorems 5.5 and 5.6 were given in Győry (1976).
@@
With the above formulation of Corollaries 5.3, 5.4 and Theorem 5.6 it was easier to point out that these are indeed consequences of Theorems 4.2 and 5.2. Further, we note that their explicit versions can be easily derived from the explicit variant Theorem 4.2* of Theorem 4.2. Finally, the corollaries can be deduced with better bounds from less general versions of Theorem 4.2, where the polynomials $f$ involved are irreducible; for such versions we refer to Győry (1976, 1998, 2000), Evertse and Győry (2017) and in fact Corollary 5.3 above.
5.4. Reformulation of Corollaries 5.3, 5.4 and Theorem 5.6 over $\mathcal{O}_{K}$ in terms of polynomial Diophantine equations over $\mathbb{Z}$
@@
Let $K$ be an algebraic number field of degree $n\geq 2$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$. Consider Corollaries 5.3, 5.4 and Theorem 5.6 for $\mathcal{O}=\mathcal{O}_{K}.$ Let $\{1,\omega_{2},\ldots,\omega_{n}\}$ be an integral basis of $K$. For $\alpha\in\mathcal{O}_{K}$ with
@@
can be regarded as a decomposable form of degree n(n1)n(n-1) in $x_{2},\ldots,x_{n}$ with coefficients in $\mathbb{Z}$, i.e., it is a product of n(n1)n(n-1) linear forms in $x_{2},\ldots,x_{n}$ with algebraic coefficients. The form $D(x_{2}\omega_{2}+\cdots+x_{n}\omega_{n})$, which was introduced by Kronecker (1882), is called discriminant form, while, for $D\neq 0$, the equation
@@
Clearly, Corollary 5.3 implies the following
@@
to the integral basis $\{1,\omega_{2},\ldots,\omega_{n}\}$ of $K$ there corresponds a decomposable form $I(X_{2},\ldots,X_{n})$ of degree n(n1)/2n(n-1)/2 in n1n-1 variables with coefficients in $\mathbb{Z}$ such that for $\alpha\in\mathcal{O}_{K}$
@@
We note that the equations (5.6) and (5.8) are related by the first identity in (5.5).
In view of (5.7), the finiteness assertion of Corollary 5.4 for $\mathcal{O}_{K}$ implies the following.
@@
In particular, for $I=1$, we get the following consequence of Theorem 5.5.
@@
Corollaries 5.7, 5.8 and 5.9 were proved in Győry (1976) with explicit upper bounds for the solutions, not only for equations (5.6), (5.8) and (5.9) but also for index form equations related to indices with respect to arbitrary orders $\mathcal{O}$ of $K$; see also Győry (2000) and Evertse and Győry (2017).